# MLM: Learning Multi-task Loco-Manipulation Whole-Body Control for Quadruped Robot with Arm

### v1

@@ III-B Network
The training framework consists of a policy network and a Critic network. The policy network acts as the Actor and controller, receiving only observable states as input.
Trajectory-Velocity Prediction Policy Network: We propose a Trajectory-Velocity Prediction Policy Network (TVP), which predicts unobservable target trajectories and body velocity, integrates additional proprioception, and outputs joint actions. A trajectory encoder $E_{\theta_{1}}$ extracts features from the historical pose frames. Then, these features are fed into a decoder $D_{\theta_{2}}$ to predict the subsequent trajectory sequence, as shown in (2). This prediction is essential, particularly in scenarios such as teleoperation, where only historical trajectory records are available. Alternatively, when deploying the strategy that utilize DP, the future trajectory can be provided directly, and the TVP will be skipped.
@@
We introduce a Normalized Autoencoder (NAE) to formulate the structure and loss function for the trajectory predictor. The reconstruction loss measures how well the NAE reconstructs the future trajectory from the latent embedding $\hat{z}_{t}^{tcp}$:
@@
Specifically, the MSE is used to minimize the discrepancy between the predicted future trajectory and the ground truth. Importantly, we apply the tanhtanh activation function to the encoded latent embedding for normalization, constraining them within a limited range. This helps prevent outliers that lie outside the expected distribution, thus aiding in stabilizing the training process. Additionally, we update the trajectory predictor, constructed by NAE, concurrently with other networks. On one hand, this facilitates accurate prediction of future trajectories; on the other hand, it generates useful latent embedding for use by the low-level network, thereby promoting the learning of the desired actions.
@@
The estimation is important because loco-manipulation tasks involve both long-range movement and short-range local operations. The estimated linear velocity of the body effectively captures the distinction between these two types of tasks, enabling more accurate execution of these different loco-manipulation tasks. The encoded embedding, along with other proprioceptive data and previous actions, is fed into a low level network, which then outputs the mean vector $\boldsymbol{\mu}_{t}^{b}\in{\mathbb{R}^{18}}$ of a Gaussian distribution $\boldsymbol{a}_{t}^{b}\sim\mathcal{N}\left(\boldsymbol{\mu}_{t}^{b},\boldsymbol{\sigma}^{b}\right)$, where $\boldsymbol{\sigma}^{b}\in{\mathbb{R}^{18}}$ represents the variance of the action, optimized by proximal policy optimization (PPO).
Critic Network: The Critic network comprises a trajectory encoder $E_{\psi_{1}}$ and a height encoder $E_{\psi_{2}}$, which perform dimensionality reduction and representation. $\psi_{1}$ and $\psi_{2}$ represent the Critic network parameters. These encoded features, combined with proprioceptive observations, previous actions and privilege information, are used to compute the state value function $V_{t}$, assessing the expected reward for the current state:
@@ III-C Trajectory Generation and Sampling
Trajectory generation encompasses both the trajectories learned during training and those tracked during deployment. The following details the generation and sampling strategies for these trajectories.
Trajectory Library in Training: We incorporate a trajectory library in whole-body policy training, which contains real-world TCP pose data from multiple manipulation tasks. During training, the initial points of these trajectories are set as the end-effector poses following episodes reset, preventing low-quality samples in the early stages.
We select trajectories from several typical manipulation tasks for learning [11]. Each task includes 200 trajectories, such as ‘clean table’, ‘pour coke’, ‘open container’, ‘pick bear’, ‘pick cup’, and ‘unplug charger’. More trajectories can be added to expand multi-task learning. We also incorporate ‘pushing’ data from [16]. These tasks cover repetitive motions, pick-and-place actions, rotational opening and closing, and long-distance movements, making them representative of common real-world operational tasks. The data are stored as separate pickle files for training. Furthermore, we find that with the guidance of the TCP trajectory tracking reward $r^{g}$, the robot can generalize to tracking other unseen trajectories to some extent.
@@
where $\bar{r}_{n}^{g}$ denotes the average tracking reward for each task trajectory. To prevent forgetting, this mechanism retains a minimum probability for each trajectory.
Meanwhile, each task trajectory adopts its own reward coefficient factor. Specifically, after repeated training and adjustments, a predefined set of average pose error thresholds and task reward factors is established as $\left\{\left(T_{k,n},\sigma_{k,n}\right)\right\}$. The average pose error for each task is calculated as follows, including Euclidean distance in displacement space and angular error in rotation space.
@@
The factor is adaptively updated based on changes in the reward values for each task trajectory. As the training progresses, this factor increases the tracking requirements for each task. The task-specific reward factor focuses specifically on the tracking performance for each individual task trajectory.
In Deployment: During deployment, the target trajectory is provided either by the pre-trained DP or through human teleoperation using the handheld gripper.
The target trajectory output by the DP is applied to an automated quadruped robot loco-manipulation task. The policy input includes multimodal data such as the arm’s proprioception and corresponding fisheye images at each time step, with the TCP trajectory defining the action space. We train the DP using the ViT-based CLIP visual encoder (ViT Base Patch 16, input 224×224), following the methods in [32] and [33].
Using the handheld gripper for teleoperation is another option that allows humans to directly control the whole-body behaviors of the quadruped robot with arm. As mentioned earlier, the whole-body policy can generalize to track previously unseen trajectories within a reasonable operational range. As a result, we developed a teleoperation scheme to control the quadruped robot, expanding its ability to perform whole-body behaviors.
The Realsense T265 on the handheld device records pose data relative to its origin at startup. We set the robot’s initial TCP point after startup as the origin of its trajectory, ensuring alignment between the initial points of the handheld device and the robot’s TCP. As the operator moves the device, the recorded historical pose data is fed into the policy for tracking, enabling whole-body behaviors under teleoperation.
@@ IV-A System Setup
During whole-body policy training, we leverage Isaac Gym to collect data in parallel from 4,096 agents. The policy updates every 50 simulation steps, and the entire training–runs on a single NVIDIA RTX 3090 Ti GPU–takes roughly eight hours to complete 10,000 iterations. The policy controller operates at $50\text{\,}\mathrm{H}\mathrm{z}$. We implement asynchronous communication for the DP or teleoperation, whole-body policy, and robot control through multi-thread control, as shown in Figure 3.
Figure 3: System Deployment Architecture: We developed a multi-computer, multi-threaded system for deployment. The host computer runs the DP ($10\text{\,}\mathrm{H}\mathrm{z}$) to generate or capture the handheld gripper’s trajectory and commands ($100\text{\,}\mathrm{H}\mathrm{z}$) using the Realsense T265 and GoPro. The frequency affects task execution speed. This reference trajectory is sent to the lower computer via Ethernet. The Jetson Orin NX, mounted on the robot, handles multi-threaded tasks, including receiving robot status ($500\text{\,}\mathrm{H}\mathrm{z}$), whole-body policy inference ($50\text{\,}\mathrm{H}\mathrm{z}$), and sending control commands ($200\text{\,}\mathrm{H}\mathrm{z}$).
@@
The controller is deployed on a 12-DoF quadruped robot, Uniree Go2, which is equipped with a 6-DoF Airbot Play robotic arm. The arm’s end effector is fitted with a single DoF parallel gripper [11], as shown in Figure 3. The robot total weights $19.9\text{\,}\mathrm{kg}$. Both the arm and the quadruped robot are powered by a battery on Go2. Inference for the whole-body policy runs onboard the Go2’s NVIDIA Jetson Orin NX, while the DP in automated mode is inferred on a PC equipped with an NVIDIA RTX 3090Ti, with the target trajectory sent via Ethernet. In teleoperation mode, a handheld gripper transmits the trajectory and gripper commands over the same link.
@@
During deployment, it is necessary to obtain the target pose and the current TCP pose in the world coordinate system. We install a Livox Mid-360 Lidar on the robot’s body and deploy an odometry [34] to build the world coordinate system, with the initial position of the Lidar serving as the origin. During system operation, the pose of the TCP is calculated through forward kinematics (FK) and transformed into the world coordinate system. The target pose is processed similarly.
IV-B Ablation Studies
@@
We demonstrate the necessity and effectiveness of the proposed method through a series of ablation experiments conducted in simulations. For experiments with the same setup, we executed 20 trials with different seeds to record the data.
Policy network wo TVP: Using the single MLP (sMLP) network-based Actor from [16] as a baseline, we validate the necessity of our TVP policy network for control. In the absence of external perception, quadruped robot movement relies on proprioceptive information to sense the environment. Our training results indicate that, when the component that estimates state using historical proprioceptive information is removed and a single MLP maps the history proprioceptive observation to actions, the task-tracking reward for loco-manipulation tasks rises more slowly during training and converges at a lower level, as shown in Figure 4. Evaluation further shows that this policy causes larger joint torque and position fluctuations, resulting in more pronounced jitter in joint motions, as shown in Figure 5(a) and (b).
@@
Figure 4: The trajectory tracking reward (left) and total reward (right) curve during the training of three policy networks: TVP (Ours), TVP without NAE, and single MLP.
@@
TVP policy network wo NAE: When the future target trajectory is unobservable (teleoperation), we carried out an ablation experiment to assess the performance conferred by the NAE, which predicts forthcoming trajectories from past trajectory data. In this setting, only the historical and current trajectories are encoded, and the resulting latent embedding is fed directly into the low-level network to generate action outputs. Because the policy lacks information about the trajectory it subsequently track, its tracking performance deteriorates: during training, the tracking reward remains lower than that achieved by the policy equipped with the NAE-based prediction mechanism, as illustrated in Figure 4. Figure 5(c) shows the MSE of the future trajectory predicted by the NAE decoder and its ground truth. The reconstruction error is represented by the position and the first two rows of pose matrix for four frames and plotted as a heatmap. The results demonstrate that the reconstruction error is minimal, indicating that the latent embedding of the NAE contains information essential for predicting the future trajectory.
@@
Figure 5: Policy network ablation analysis: (a) Schematic of the ‘unplug charger’ and ‘pushing’ action sequences in simulation. (b) Arm joint torque and position curves for TVP and single MLP networks during the tasks, with sMLP showing more jitter. (c) MSE heatmap of future trajectory reconstruction by TVP’s NAE for both tasks.
@@
Table III shows the policy evaluation for seven tasks. The position and orientation errors of the end-effector’s trajectory tracking were recorded after 800 time steps. Position error is measured using Euclidean distance (cm), while orientation error refers to the minimum rotational angle between the end-effector and the target trajectory’s orientation (radian). The average error over all time steps was computed to assess overall tracking performance. Results show that the TVP, with its future trajectory prediction mechanism, outperforms other policies, while the single MLP-based policy performs the worst due to its limited information extraction and representation ability.
@@
Adaptive trajectory sampling: In this experiment, we set all robots to uniformly sample (Uni-sample) all task trajectories throughout the training process in order to assess the necessity of our adaptive trajectory sampling strategy. We record the reward curves for two typical tasks during training, as shown in Figure 6. The results show that uniform sampling cannot effectively balance task performance. Initially, the reward curves for tasks with different pose variations are close, but over time, the rewards diverge, with tasks requiring more pose variation converging to lower rewards due to insufficient attention (i.e., higher sampling rates). In contrast, adaptive sampling maintains similar rewards across tasks, balancing tracking performance. Table III indicates that tasks with more frequent trajectory changes (e.g., ”open container”) see a significant decline in tracking accuracy.
@@
TABLE III: Simulation ablation study: mean 6D tracking error for each task, and the variance of those mean errors across all tasks. Metrics: Euclidean distance / rotation-matrix angular difference. (cm/radian)
@@
TVP wo NAE 1.78/0.11 1.49/0.07 1.30/0.13 1.26/0.07 1.43/0.05 1.31/0.10 1.78/0.22 -
sMLP (baseline) 2.23/0.12 2.02/0.14 2.31/0.37 2.24/0.14 2.20/0.13 2.17/0.22 3.13/0.54 -

### v2

@@ III-B Network
Trajectory-velocity prediction policy network: We propose a Trajectory-Velocity Prediction policy network (TVP) consisting of an NAE, a memory encoder, an estimator and a low-level network. The historical trajectories are processed by the NAE to first obtain the NAE embeddings and then predict future poses. The historical proprioceptive state sequence is fed into the memory encoder (ME) to produce the ME embeddings. The ME embeddings are used as inputs to the estimator, which predicts the body’s linear velocity. The predicted future poses, the NAE embeddings, the predicted linear velocity, the ME embeddings, the current proprioceptive state and the last action are fed into a low-level network to generate the joint action.
A trajectory encoder $E_{\theta_{1}}$ extracts features from the historical poses. Then, these features are fed into a decoder $D_{\theta_{2}}$ to predict the subsequent trajectory sequence, as shown in (2). This prediction is essential, particularly in teleoperation, where only historical trajectory records are available. Alternatively, when deploying the strategy that utilize DP, the future trajectory can be provided directly, and the TVP will be skipped.
@@
We introduce a normalized autoencoder (NAE) to formulate the structure and loss function for the trajectory predictor. The reconstruction loss measures how well the NAE reconstructs the future trajectory from the embeddings $\hat{z}_{t}^{tcp}$:
@@
Specifically, the MSE is used to minimize the discrepancy between the predicted future trajectory and the ground truth. Importantly, we apply the tanhtanh activation function to the encoded embeddings for normalization, constraining them within a limited range. This helps prevent outliers that lie outside the expected distribution, thus aiding in stabilizing the training process. Additionally, we update the trajectory predictor, concurrently with other networks. This both facilitates accurate prediction of future trajectories and provides useful embeddings for the low-level network, thereby promoting the learning of the desired action.
@@
The estimation is important because loco-manipulation tasks involve both long-range movement and short-range local operations. The estimated linear velocity of the body effectively captures the distinction between these two types of tasks, enabling more accurate execution of these different loco-manipulation tasks. The encoded embeddings and predicted values, together with current proprioceptive state and the last action, are fed into a low level network, which then outputs the mean vector $\boldsymbol{\mu}_{t}^{b}\in{\mathbb{R}^{18}}$ of a Gaussian distribution $\boldsymbol{a}_{t}^{b}\sim\mathcal{N}\left(\boldsymbol{\mu}_{t}^{b},\boldsymbol{\sigma}^{b}\right)$, where $\boldsymbol{\sigma}^{b}\in{\mathbb{R}^{18}}$ represents the variance of the action, optimized by proximal policy optimization (PPO).
Critic network: The critic network comprises a trajectory encoder $E_{\psi_{1}}$ and a height encoder $E_{\psi_{2}}$, which perform dimensionality reduction and representation. $\psi_{1}$ and $\psi_{2}$ represent the critic network parameters. These encoded features, combined with proprioceptive observations, last action and privilege information are used to compute the state value function $V_{t}$, assessing the expected reward for the current state:
@@ III-C Trajectory Generation and Sampling
Trajectory generation encompasses both training and deployment, with the subsequent sections detailing the corresponding generation and sampling strategies.
Trajectory library in training: We introduce a trajectory library containing real-world TCP poses from multiple manipulation tasks. During training, the initial points of these trajectories are set as the end-effector poses following episodes reset, preventing low-quality samples in the early stages.
We select trajectories from several typical tasks for learning, such as ‘clean table’, ‘pour coke’, ‘open container’, ‘pick bear’, ‘pick cup’, and ‘unplug charger’ [Zhaxizhuoma2025FastUMI]. Each task includes 200 trajectories. More trajectories can be added to expand multi-task learning. We also incorporate ‘pushing’ from [Huy2024UMIonLegs]. These tasks cover repetitive motions, pick-and-place, rotational opening and closing, and long-distance movements, making them representative of common real-world manipulation tasks. The data is stored as separate pickle files. Furthermore, guided by the tracking reward, the robot demonstrates a certain ability to generalize to previously unseen trajectories.
@@
where $\bar{r}_{n}^{g}$ denotes the average tracking reward for each task trajectory. $\lambda_{n}$ is the reward performance threshold for task $n$. By tuning the values of $\lambda_{n}$, we can adjust the sampling probability curves of the reference trajectories of different tasks, so that we can achieve similar reward levels for different tasks. $\lambda_{n}$ is determined by the user according to reward curves of all tasks after policy training. To prevent forgetting, this mechanism retains a minimum probability for each trajectory.
Meanwhile, each task trajectory adopts its own reward coefficient factor. Specifically, after repeated adjustments, a predefined set of average pose error thresholds and task reward factors is established as $\left\{\left(T_{k,n},\sigma_{k,n}\right)\right\}$. The average pose error for each task is calculated as follows, including Euclidean distance and angular error.
@@
The factor is adaptively updated based on changes in the reward values for each task trajectory. As the training progresses, this factor increases the tracking requirements for each task.
In deployment: During deployment, the target trajectory can be provided either by teleoperation or by pre-trained DP.
Using the handheld gripper for teleoperation allows humans to directly control the quadruped robot with an arm, and expands the robot’s ability to perform whole-body behaviors. As mentioned earlier, the policy can generalize to track previously unseen trajectories within a reasonable range. The handheld device uses the RealSense T265 to record poses. At startup, we set the arm’s TCP as the trajectory origin, aligning it with the T265’s initial point. As the operator moves, the recorded poses are fed into the policy for tracking. Meanwhile, GoPro-based marker distance estimation directly triggers gripper commands.
The target trajectory output by the DP is applied to an automated loco-manipulation task. In our case, the DP is a pre-trained module using the training approach in [Chi2023Diffusion, Zhaxizhuoma2025FastUMI]. In the DP training, a ViT-based CLIP visual encoder [Radford2021Learning] is employed with ViT Base/16 architecture and an image resolution of 224×224. The policy input includes the arm’s proprioception and fisheye images at each time step, with the target trajectory defining the output.
@@ IV-A System Setup
We use Isaac Gym to collect data from 4,096 agents for whole-body policy training. The policy is updated every 50 simulation steps and trained on an NVIDIA RTX 3090Ti, requiring about 8 hours for 10,000 iterations.
@@
As shown in Figure 3, the controller is deployed on a 12-DoF Uniree Go2 quadruped robot with a 6-DoF Airbot Play robotic arm. The arm’s end-effector is fitted with a single DoF parallel gripper [Zhaxizhuoma2025FastUMI]. The robot total weights $19.9\text{\,}\mathrm{kg}$. Both the arm and the quadruped robot are powered by a battery on Go2. We implement asynchronous communication for our system through multi-thread. The whole-body policy runs onboard the Go2’s NVIDIA Jetson Orin NX at $50\text{\,}\mathrm{H}\mathrm{z}$. Inference of DP and the handheld gripper run on a PC with an NVIDIA RTX 3090 Ti that transmits the target trajectory via Ethernet.
@@
During deployment, it is necessary to obtain the target and end-effector poses in the world frame. A Livox Mid-360 LiDAR is mounted on Go2, and an odometry [Chen2023Direct] is used to build the world frame, with the LiDAR’s initial position as the origin. During inference, the poses are calculated through forward kinematics and transformed into the world frame.
Figure 3: A multi-computer, multi-thread system deployment architecture. The host computer captures the handheld gripper’s trajectory ($100\text{\,}\mathrm{H}\mathrm{z}$) or runs the DP ($10\text{\,}\mathrm{H}\mathrm{z}$). The trajectory is sent to the lower computer via Ethernet. The Jetson Orin NX, mounted on the robot, handles multi-thread tasks, including receiving robot states ($500\text{\,}\mathrm{H}\mathrm{z}$), whole-body policy inference ($50\text{\,}\mathrm{H}\mathrm{z}$), and sending control commands ($200\text{\,}\mathrm{H}\mathrm{z}$).
IV-B Simulation Studies
@@
We demonstrate the necessity and effectiveness of the proposed method through simulation experiments. For experiments with the same setup, we executed 20 trials with different seeds.
Policy network wo TVP: Using a single MLP (sMLP) actor from the state-of-the-art work [Huy2024UMIonLegs] as a representative baseline, which has demonstrated good performance on pushing and cup-picking tasks, we highlight the necessity of our TVP policy network for control. This baseline policy does not predict the base velocity. The historical and current trajectories, the current proprioceptive state and the last action are fed into a single MLP network to generate the joint action. Our results show that the baseline policy reward exhibits slower growth and achieves a lower level of convergence, as shown in Figure 4. Evaluation further shows that this baseline policy causes larger joint torque and position fluctuations, leading to more noticeable joint jitter, as shown in Figure 5(a) and (b). To further quantify the jitter differences, we introduce a mean squared derivative (MSD) of arm joint torques over a time interval, $MSD=\frac{1}{(t_{1}-t_{0})/dt}{\textstyle\sum_{t=t_{0}}^{t=t_{1}-dt}(\tau_{t+dt}-\tau_{t})^{2}}$. In our case, the time interval is from $t_{0}=5s$ to $t_{1}=10s$. The results of ‘unplug charger’ in Table III show that the TVP can achieve smoother torque control, as it has the lower MSD than sMLP.
@@
Figure 4: The trajectory tracking reward (left) and total reward (right) curve during the training of four policy networks: TVP (Ours), TVP wo NAE, single MLP, and DeepWBC.
TABLE III: MSD of six joint torques.
Method Joint 1 Joint 2 Joint 3 Joint 4 Joint 5 Joint 6
TVP 0.0254 0.2660 0.2144 0.0021 0.0026 0.0029
sMLP 0.0515 0.2867 0.2500 0.0098 0.0038 0.0039
@@
TVP policy network wo NAE: In teleoperation, where the future trajectory is unobservable, we conducted an ablation study to evaluate the NAE, which predicts future trajectories from past data. In this ablation setting, the historical proprioceptive state sequence is fed into the ME to produce the ME embeddings. The ME embeddings are used as inputs to the estimator, which predicts the body’s linear velocity. The predicted linear velocity, the ME embeddings, the historical and current trajectories, the current proprioceptive state, and the last action are fed into a low-level network to generate the joint action. As shown in Figure 4, without trajectory prediction, the policy lacks guidance on “where to go next”, leading to degraded tracking performance: the tracking reward remains lower than that of the NAE-based policy. Figure 5(c) shows the MSE between the NAE-predicted trajectory and the ground truth. The reconstruction errors of position and the first two pose-matrix rows over four frames are visualized as a heatmap. The results demonstrate that the reconstruction error is minimal, indicating that the NAE embeddings contain information essential for predicting the future trajectory.
@@
Figure 5: Policy network ablation analysis: (a) The ‘unplug charger’ and ‘pushing’ trajectories in simulation. (b) Arm joint torque and position curves for TVP and sMLP during the tasks, with sMLP showing more jitter. (c) MSE heatmap of future trajectory reconstruction by NAE for both tasks.
@@
Table IV shows the policy evaluation for seven tasks. The position and orientation errors of the end-effector’s trajectory tracking were recorded over 800 time steps. Position error is measured using Euclidean distance (cm), while orientation error refers to the minimum rotational angle (radian). The average error over all time steps was computed to assess overall tracking performance. Results show that the TVP, with its future trajectory prediction mechanism, outperforms other policies. The sMLP policy performs the worst due to its limited information extraction and representation ability. The error sources could include the dynamics and noise of the command trajectory, the dynamics of the robots, and the policy network.
Policy wo NAE, task-frame target and trajectory input: We also add a comparison with DeepWBC [Fu2023DeepWholeBody]. DeepWBC is obtained as follows. We remove the NAE that predicts future trajectories and discard trajectories as the policy targets, relying instead on a single target point. The end-effector pose is defined in the body frame. Through the above setting, we can obtain a method similar to DeepWBC. The experimental results for each task are presented in Table IV. Because the target is defined in the world frame rather than the body frame, our policy is not sensitive to body motion. Because our method has continuous trajectories as reference targets and has a prediction mechanism, our policy’s tracking performance is better than that of DeepWBC.
@@
Adaptive trajectory sampling: To assess the necessity of adaptive trajectory sampling mechanism, we set all robots to uniformly sample (Uni-sample) task trajectories throughout the training. We record the reward curves for two typical tasks during training, as shown in Figure 6. The results show that uniform sampling cannot effectively balance task performance. Initially, reward curves for tasks with different pose variations are close. Over time, the rewards diverge. Tasks with more pose variation converge to lower reward due to insufficient attention. In contrast, adaptive sampling maintains similar rewards across tasks, balancing tracking performance. Table IV indicates that tasks with frequent trajectory changes (e.g., ”open container”) see a significant decline in tracking accuracy.
@@
Moreover, the TVP (Ours) and the TVP wo NAE have lower variances than the baselines (sMLP and DeepWBC). Although the TVP (Ours) has a similar variance to the TVP wo NAE, the TVP (Ours) has the lowest mean tracking errors across all tasks.
Impact of prediction errors: To evaluate the impact of prediction errors on policy performance, we inject Gaussian noise with increasing magnitudes into both the position (1 cm, 2 cm, 10 cm) and orientation (0.05 rad, 0.1 rad, 0.5 rad) components of the predicted trajectory. The experimental results are reported in Table IV. As the prediction errors grow, the tracking errors of the loco-manipulation also rise, but the increase remains acceptable, and the tasks can still be completed successfully. In extreme cases, where the predictor outputs unreasonable future steps (e.g., Gaussian noise with a mean of 10 cm and 0.5 rad), the loco-manipulation still functions. The reason is that our method does not rely solely on the predicted trajectory but also on the historical ground-truth trajectory, which provides good robustness against prediction errors.
TABLE IV: Simulation ablation study: mean 6D tracking errors for each task, and the variance of those mean errors across all tasks. Metrics: Euclidean distance / rotation-matrix angular difference. (cm / radian)
@@
TVP wo NAE 1.78/0.11 1.49/0.07 1.30/0.13 1.26/0.07 1.43/0.05 1.31/0.10 1.78/0.22 0.049/0.0032
sMLP [Huy2024UMIonLegs] 2.23/0.12 2.02/0.14 2.31/0.37 2.24/0.14 2.20/0.13 2.17/0.22 3.13/0.54 0.133/0.0256
DeepWBC [Fu2023DeepWholeBody] 8.32/0.51 8.05/0.21 6.83/0.49 7.35/0.21 7.09/0.23 9.11/0.30 7.22/0.61 0.663/0.0279
Noise ($N(1,0.2^{2})/N(0.05,0.01^{2})$) 1.43/0.10 1.23/0.08 0.92/0.11 0.79/0.05 0.91/0.06 1.08/0.09 1.26/0.16 0.053/0.0013
Noise ($N(2,0.4^{2})/N(0.1,0.02^{2})$) 1.59/0.11 1.32/0.08 0.99/0.12 0.88/0.05 1.07/0.06 1.22/0.10 1.33/0.17 0.058/0.0016
Noise ($N(10,2^{2})/N(0.5,0.1^{2})$) 2.19/0.13 1.51/0.11 1.28/0.15 1.50/0.07 1.40/0.09 1.39/0.11 1.79/0.25 0.098/0.0035